### v1

# On the interpretation of quantum theory as games between physicists and nature played in Minkowski spacetime

Abstract
In 2019, we introduced games in Minkowski spacetime as a generalization of game theory to special relativity that subsumes games in normal form (spacelike separation) and games in extensive form (timelike separation). Many concepts including Nash equilibria naturally extend to spacetime games. We also emphasized the importance of these games to model quantum experiments such as Bell experiments and more generally any adaptive measurements. Subsequent work suggested to formalize a special case of such games in terms of strategy presheaves. In the case that measurements have a unique causal bridge and if a natural cover is taken, we show that the two frameworks are isomorphic to each other and provide complementary perspectives. Spacetime games provide a visual and intuitive framework that also captures the distinction between joint experiments and either-or experiments, so that they are rich enough in their causal structure to imply a natural cover for the corresponding causal contextuality scenario. Based on this observation, we suggest to define the strategy presheaf directly based on the pure strategies (and restrictions thereof) of the spacetime game, and we show that the sheaf property obtains for the games at hand. The argument is rather simple and similar to event sheaves for the flat case. Finally, we explain how, in the other direction, the failure of the sheaf property on strategy distribution presheaves is consistent with our previous argument that Nash game theory is not compatible with quantum physics. This shows that the insights of the two frameworks, taken together, can contribute positively to the advancement of the field of quantum foundations.
1 Introduction and background
Fourny (2019a) and Fourny (2020) introduced a new class of games called spacetime games with perfect information, in order to extend the semantics of game theory to special relativity. In spacetime games, the agents make decisions at various positions in Minkowski spacetime. Spacetime games can be seen as the least common denominator of strategic games on the one hand (spacelike separation), and games in extensive form with perfect information on the other hand (timelike separation).
It is explained in the same reports how these games can be used to model quantum experiments as a game played between experimenters and nature. An example of game corresponding to the Bell experiment is shown in Figure 1 (Alice and Bob are the experimenters and Alfred, which we like to visualize as a cat, represents nature):
Figure 1: A spacetime game with perfect information that represents a Bell experiment with two spacelike experimenters.
A canonical injection of the class of spacetime games with perfect information into the class of games in extensive form with imperfect information is also given. Figure 2 shows the game in extensive form with imperfect information (dashed lines group nodes into information sets, payoffs are omitted) that corresponds to the game of Figure 1. The information sets of the extensive form in Figure 2 correspond exactly to those of the spacetime game in Figure 1 (which are singletons, one for each node, because the spacetime game has perfect information). This is always the case for any spacetime games and their extensive form.
Figure 2: The game in extensive form with imperfect information that correponds to the spacetime shown on Figure 1. Alice and Bob have each one information set for their choice of measurement setting. Alfred has four information sets X, Y, W and Z, one for each measurement.
Subsequent work by Abramsky et al. (2024) suggests to model a subclass of such games, which they call causal contextuality scenarios, as strategy presheaves. In this paper, we show that when causal bridges are unique and when natural covers are taken, causal contextuality scenarios are isomorphic to a subclass of spacetime games, which we call alternating spacetime games. We also show that a natural definition of the strategy presheaf emerges, and that the sheaf property holds. We suggest that the isomorphism can be extended to non-unique causal bridges and to a broader subclass of spacetime games, which we leave open for future work.
1.1 Spacetime games
For the purpose of defining the isomorphism, we now give a compact, formal definition of spacetime games. In this definition, we do not require perfect information. Imperfect information has the meaning proposed by Kuhn (1950) and extends naturally to spacetime games without affecting the injection into the class of games in extensive form with imperfect information. [^1] This is because merging nodes into information sets in the spacetime game translates directly into merging the corresponding information sets in the extensive form.
Definition 1 (Spacetime game).
A spacetime game $\mathcal{G}$ is a tuple $(\mathcal{N},\mathcal{R},\mathcal{P},\mathcal{A},\rho,\chi,$ $\sigma,\mathcal{I},\mathcal{Z},u)$ where
  • $(\mathcal{N},\mathcal{R})$ is a directed acyclic graph of decision nodes, $\mathcal{N}$ being the nodes and $\mathcal{R}$ being the edges.
  • $\mathcal{P}$ is a set of players and $\mathcal{A}$ is a set of actions (in our context, measurement settings and measurement outcomes)
  • Nodes are labeled with players: $\rho\in\mathcal{P}^{\mathcal{N}}$
  • Nodes are labeled with sets of actions: $\chi\in\mathcal{P}(\mathcal{A})^{\mathcal{N}}$
  • Edges are labeled with actions: $\sigma\in\mathcal{A}^{\mathcal{R}}$ labels each edge $(N,M)$ with an action $\sigma(N,M)\in\chi(N)$.
  • a partition of $\mathcal{N}$ into information sets: a set $\mathcal{I}$ of subsets of $\mathcal{N}$ that form a partition, in a way that is compatible with $\rho$ and $\chi$, meaning that
    $\forall i\in\mathcal{I},\forall N,M\in i,\rho(N)=\rho(M)\wedge\chi(N)=\chi(M)$
  • $\mathcal{Z}\subseteq(\mathcal{A}\cup\{\bot\})^{\mathcal{I}}$ is a set of outcomes, which are partial functions from information sets to actions. The outcomes are the complete histories mapping nodes to actions. They must obey some constraints implied by the game structure, and are defined further down in Section 1.3. The domain of definition of an outcome is the set of information sets not mapped to $\bot$, which is the symbol used for unassigned actions.
  • $u=(u_{i})_{i\in P}$ is a family of utility functions mapping outcomes to payoffs, with one payoff for each outcome and player.
The payoffs associated with the outcomes are part of the semantics of a game, but are not the focus of the present paper, which focuses on the structure of the games. They are also not part of any causal contextuality scenario and ignored in the mappings.
Figure 3 shows an example of spacetime game with imperfect information (the dashed lines have the same meaning as on Figure 2) and Figure 4 shows the corresponding extensive form with imperfect information.
Figure 3: A spacetime game with imperfect information that represents a contextuality experiment with two settings to pick out of three. Bob has one information set with three actions (the three contexts XY, XZ and YZ). Alfred has 3 information sets X, Y and Z for each measurement, with two actions each, 0 and 1.
Figure 4: The game in extensive form with imperfect information that corresponds to the spacetime game shown on Figure 3. It has exactly the same information sets: XY, YZ and XZ for Bob, and X, Y and Z for Alfred.
Figure 5 shows an example of spacetime game with perfect information that has a causal bridge and the corresponding extensive form is shown in Figure 6. Spacetime games can be much more complex and mix spacelike-separation as in Figure 1, timelike-separation as in Figure 5, and imperfect information for explicit contexts as shown in Figure 8.
Figure 5: A spacetime game with imperfect information that shows an example of causal bridge: Bob only carries out his experiment if Alice measured Y and obtained an outcome of 0.
Figure 6: The game in extensive form with imperfect information that corresponds to the spacetime game shown on Figure 5. It happens to have the same tree shape, because there is no spacelike separation
1.2 Notations
In the remainder of this paper, whenever we refer to a game $\mathcal{G}$, we will also implicitly use the above notations ($\mathcal{N},\mathcal{R}$, $\mathcal{P}$, $\mathcal{A}$, $\rho$, $\chi$, $\sigma$, $\mathcal{I}$) without explicitly defining them every time. This is because we only ever consider one game at a time, and it eases the read.
We write $N\smile M$ as a convenient shorthand for $(N,M)\in\mathcal{R}$.
We define for any node N
$\textsf{precedessors}(N)=\{M\in\mathcal{N}|M\smile N\}$
$\textsf{successors}(N)=\{N\in\mathcal{N}|N\smile M\}$
We also define for any information set $i$
$\textsf{precedessors}(i)=\bigcup_{N\in i}\textsf{predecessors}(N)$
$\textsf{successors}(i)=\bigcup_{N\in i}\textsf{successors}(i)$
We also use a slight abuse of notation whenever all nodes in the same information set share a common property: for example, $\rho(i)$ and $\chi(i)$ where $i$ is an information set. We also use, for convenience, a function $\iota\in\mathcal{I}^{\mathcal{N}}$ that maps each node to the information set it belongs to.
Conversely, we will write $z(N)$ for $z(\iota(N))$ where $z$ is an outcome (complete history) and $N$ is a node.
We also write $\textsf{parent}(N)$ for the at most unique $P\in\mathcal{N}$ such that $P\smile N$ when it is known from the context that there is a unique predecessor; if it is known that there is none, then this is the empty set. The abuse of notation for information sets then translates to $\textsf{parent}(i)$ where $i$ is an information set for which it is known that there is a unique common predecessor to all its nodes. Finally, when we write the statement $\sigma(N,M)=a$ for some $a$, it implicitly implies $N\smile M$, which allows more concise formulas.
1.3 Histories
We now give a compact definition of the histories of a spacetime game:
Definition 2 (Histories).
Given a spacetime game in extensive form $\mathcal{G}$, a history is an assignment from $\mathcal{I}$ to $\mathcal{A}\cup\{\bot\}$ if it only assigns an action to activated information sets. An information set is activated if it has a node such that the label of each one of its incoming edges is equal to the action assigned (by that same history) to the information set containing the source node of that edge:
$\forall i\in\mathcal{I},(z(i)\neq\bot\implies\exists N\in i,\forall M\in\textsf{predecessors}(N),z(M)=\sigma(M,N))$
The set of all histories is denoted $\mathcal{H}$.
Fourny (2020) uses the term “contingency coordinates” of a node $N$ to denote the minimal history not containing $N$ that coincides with any history in which $N$ is assigned an action (i.e., not $\bot$). In this paper, we use the term “causal bridge” (Fourny et al., 2014) [^2] The idea of a decision happening only if a certain causal past is secured, and otherwise not, can be traced back to (Dupuy, 1992) and (Dupuy, 2000), and we retain his terminology “causal bridge”. of a node $N$ to specifically designates the predecessors of $N$ and the labels on the edges from the predecessors to $N$.
A history is complete if the implication becomes an equivalence. In a valid game, the set of outcomes is the set of complete histories.
Definition 3 (Complete histories, and validity of a spacetime game).
A spacetime game in extensive form $\mathcal{G}$ if valid if the set of outcomes $\mathcal{Z}$ matches the set of complete histories. A complete history is defined as a history that assigns an action to all the activated information sets.
$\forall i\in\mathcal{I},(z(i)\neq\bot\iff\exists N\in i,\forall M\in\textsf{predecessors}(N),z(M)=\sigma(M,N))$
We always assume that spacetime games are valid, as we consider validity to be part of their definition. Given a complete history (outcome) $z$, we write
$\textsf{domain}(z)=\{i\in\mathcal{I},z(i)\neq\bot\}$
Definition 4 (Clean spacetime game).
A spacetime game is clean if it additionally fulfils the following criteria:
  • It does not contain any information sets that are never assigned any value in any complete history.
    $\forall i\in\mathcal{I},\exists z\in Z,z(i)\neq\bot$
  • It does not contain any unused label.
    $\forall i\in\mathcal{I},\forall a\in\chi(i),\exists z\in Z,z(i)=a$
  • The function $\chi$ does not return unused actions at any information sets, i.e., such that no edge from that node is labeled with them.
    $\forall i\in\mathcal{I},\forall a\in\chi(i),\exists N\in i,\exists M\in\mathcal{N},\sigma(N,M)=a$
  • The set $\mathcal{A}$ does not contain any unused actions.
    $\mathcal{A}=\bigcup_{N\in\mathcal{N}}\chi(N)$
Without loss of generality, we assume the game is clean. It is always possible to remove unused information sets, to remove unused actions from sets returned by $\chi$ and from $\mathcal{A}$, to remove nodes that have an incoming edge with an unused label, and to merge nodes.
2 Results
We now formally show that, for a subclass of two-player spacetime games (satisfying an “alternating” property) we can build an equivalent causal contextuality scenario. We also show that, in the other direction, we can convert back causal contextuality scenarios to two-player spacetime games. In this paper, we focus on scenarios with unique causal bridges, which means that each measurement is enabled only by one causal past. See Section 3 for a discussion of extensions of the isomorphism (and of the corresponding isomorphic subclasses) to non-unique causal bridges, which are also supported by spacetime games.
Thus, spacetime games and causal contextuality scenarios are isomorphic for a broad class of experimental protocols, which can be either characterized as alternating spacetime games, or as causal contextuality scenarios with unique causal bridges and “good covers” (see Question 7.3 by Abramsky et al. (2024)), whereby we suggest that a cover is “good” is it corresponds to a natural cover implied by a spacetime game structure.
Complete histories under the two frameworks are isomorphic, and histories in causal contextuality scenarios correspond to those histories in spacetime games that have a domain with even cardinality.
2.1 Alternating spacetime games
We now give a characterization of a subclass of spacetime games called “alternating spacetime games”, for which we will show that there is an isomorphism with the corresponding causal contextuality scenarios.
Note that, since causal contextuality scenarios do not specify which experimenter performs which experiment, we call all our experimenters Bob for the purpose of the mapping. If a causal contextuality scenario is extended with information on the identity of the experimenters (e.g. Alice, Bob) then they can be marked accordingly in the corresponding game.
We include, in the definition of the alternating property, the interpretation of the assumed features of the game in terms of quantum-experimental protocols, to facilitate the understanding of the isomorphism that will be given further down.
Definition 5 (Alternating property).
A spacetime game $\mathcal{G}$ is an alternating game with respect to a player Alfred (typically embodying nature) if it is valid and clean, and:
  1. 1.
    It has two players (we call the other Bob).
  2. 2.
    The graph $(\mathcal{N},\mathcal{R})$ always connects a node played by Alfred to a node played by Bob, or a node played by Bob to a node not played by Alfred. In other words, the graph structure of the spacetime game is a bipartite graph if we ignore the direction of the edges.
  3. 3.
    All information sets played by Bob are singletons (“unique causal bridge” to the choice of setting).
  4. 4.
    All nodes in the same information set $i$ played by Alfred [^3] This trivially applies to Bob, too. share the same and unique parent node, which we denote $\textsf{parent}(i)$, i.e., all edges to nodes in this information set come from this parent node (“unique causal bridge” to the choice of outcome):
    $\forall i\in\mathcal{I},\exists P\in\mathcal{N},\forall N\in i,\textsf{predecessors}(N)=\{P\}$
  5. 5.
    All information sets played by Bob have at least one successor. In other words, if a context with settings is selected, they the corresponding experiments actually happen and yield a result.
    $\forall N\in\rho^{-1}(\text{Bob}),\textsf{successors}(N)\neq\{\}$
  6. 6.
    All nodes in the same information set played by Alfred [^4] This trivially applies to Bob, too. have the same outgoing edges: same labels, same destination nodes. In other words, the causal future of a measurement does not depend on the context in which it was carried out.
    $\forall i\in\mathcal{I},\forall N,M\in i,\textsf{successors}(N)=\textsf{successors}(M)$
    $\forall i\in\mathcal{I},\forall N,M\in i,\forall S\in\textsf{successors}(i),\sigma(M,S)=\sigma(N,S)$
  7. 7.
    Two distinct nodes played by Bob cannot have the same causal bridge. [^5] This amounts to saying that we assume that the measurements carried out by Bob under the same enabling circumstances are also “grouped” into a unique experiment where Bob has multiple contexts to choose from. Note that this does not cause any loss of generality, because it is always possible to “fix” a game that would not fulfil this property by merging nodes with the same contingency coordinates, building all combinations of information sets, and concatenating the labels.
    $\forall N,M\in\rho^{-1}(\text{Bob}),$
    $\textsf{predecessors}(N)=\textsf{predecessors}(M)\wedge\forall P\in\textsf{predecessors}(N),\sigma(P,N)=\sigma(P,M)$
    $\implies N=M$
  8. 8.
    Given a node $N$ played by Bob, two nodes (played by Alfred) connected to $N$ with the same label (it is a measurement context) must be in different information sets (these are measurement settings). In other words, distinct nodes for the same measurement in the same context would be superfluous.
    $\forall N\in\rho^{-1}(\text{Bob}),\forall S,T\in\textsf{successors}(N),\sigma(N,S)=\sigma(N,T)\implies\iota(S)\neq\iota(T)$
2.2 Standardization of Bob’s actions in alternating spacetime games
Without loss of generality, we can put constraints on Bob’s actions in an alternating game to facilitate their interpretation as contexts, like so:
  • Each edge from a node played by Bob to a node played by Alfred is labeled with a set (the context) of information sets (the settings) played by Alfred.
  • For any node $N$ played by Bob, there are as many outgoing edges from $N$ as there are context-setting pairs $(x,c)$ such that $x\in c$ and $c\in\chi(N)$. The edge corresponding to $(x,c)$ must go to a node $M\in x$ and must be labeled with $\sigma(N,M)=c$.
In other words, Bob’s actions are contexts, and for each context, there are edges to each measurement setting in that context, labeled with that context.
It is always possible to rename Bob’s actions such that this is the case: Given a node $N$ played by Bob and an action $c\in\chi(N)$, we rename $c$ to $\{\iota(M)|M\in\mathcal{N}\wedge\sigma(N,M)=c\}$. Because of alternating condition 6, the renaming is bijective.
We show below a correspondence table between the two frameworks.
Spacetime games Causal contextuality scenarios
Alfred the nature player
Bob the experimenter
An information set played by Alfred a measurement
A node played by Alfred a measurement and its context
An information set played by Bob a set of events, which is the left-hand-
side of an enabling relation
An action by Alfred a measurement outcome
An action by Bob a measurement context
An edge from Bob to Alfred Bob picks a measurement context and
carries out a measurement in this context.
The label is the context.
An edge from Alfred to Bob An outcome $o$ is obtained for a specific
measurement $x$, and is part of enabling
conditions $t$ for another measurement.
The label is $t(x)=o$
2.3 Causal contextuality scenarios
We recall the definition of a causal contextuality scenario as given by Abramsky et al. (2024):
Definition 6 (Causal contextuality scenario).
A causal contextuality scenario $\Gamma$ is a quadruplet $(X,O,\vdash,\mathcal{C})$ where X is a set of measurement settings, $O=(O_{x})_{x\in X}$ is a family of sets of measurement outcomes, indexed by settings, and $\vdash$ is an enabling relation between a section on a subset of $X$ (a “consistent set of events”) and settings, and $\mathcal{C}$ is a family of subsets of $X$ whose union is X.
2.4 Natural cover
Each spacetime game (not necessarily fulfilling the alternating property) has a natural cover that corresponds to all combinations of measurements that are performed in one of the complete histories:
Definition 7 (Natural cover associated with a game).
Given a spacetime game $\mathcal{G}$ with nature player Alfred, the natural cover $\mathcal{C}$ is defined as follows: For each outcome $z\in Z$, we take the subset $C_{z}=\{i\in\textsf{domain}(z)|\rho(i)=\text{Alfred}\}$ of its domain of definition with only and exactly the information sets that are played by Alfred. Then the natural cover is $\mathcal{C}=\{C_{z}|z\in Z\}$. Note that duplicates are implicitly eliminated because of the definition of a set.
For example, the natural cover of the game shown in Figure 1 and Figure 2 is $\{\{XW\},\{XZ\},\{YW\},\{YZ\}\}$.
The natural cover of the spacetime game shown in Figure 3 and Figure 4 is $\{\{XY\},\{XZ\},\{YZ\}\}$.
The natural cover of the spacetime game shown in Figure 5 and Figure 6 is $\{\{X\},\{YW\},\{YZ\}\}$.
This natural cover will be used to define the corresponding measurement scenario for any alternating game.
2.5 The mapping
Any alternating spacetime game can be associated with a causal contextuality scenario as follows. As explained in the previous tabular correspondence, the general idea of the overall mapping is that
  • Alfred’s information sets are measurements;
  • Bob’s (singleton) information sets are the left-hand sides of the enabling relations of the causal contextuality scenario;
  • Bob’s actions (which label edges from Bob to Alfred) are contexts. Alfred’s nodes within the same information set are measurements with the same measurement setting but in different contexts;
  • Alfred’s actions (which label edges to Bob) are labeled by measurement outcomes.
Definition 8 (Causal contextuality scenario associated with a spacetime game).
Given an alternating spacetime game $\mathcal{G}$ with nature player $\text{Alfred}\in\mathcal{P}$, we define the equivalent causal measurement scenario $\Sigma(\mathcal{G})=(X,O,\vdash)$ as follows:
  • $X=\{i\in\mathcal{I}|\rho(i)=\text{Alfred}\}$
  • for each $x\in X$, $O_{x}=\chi(x)$
  • for each $x\in X$, we define that $t\vdash x$ where
    $t=\{(\iota(M),\sigma(M,\textsf{parent}(x)))|M\smile\textsf{parent}(x)\}$
    Note that when $\textsf{parent}(x)=\{\}$, this amounts to {}$\{\}\vdash x$. Also note that it means that $\sigma(M,\textsf{parent}(x))=t(y)$ for any $M\in y$.
  • the cover $\mathcal{C}$ is the natural cover of $\mathcal{G}$.
We use the abbreviation $\textsf{enabled}(t)=\{x|t\vdash x\}$, in particular for restricting the cover $\mathcal{C}$ to $\mathcal{C}\cap\textsf{enabled}(t)$ which we define as
$\mathcal{C}\cap\textsf{enabled}(t)=\{C\cap\textsf{enabled}(t)|C\in\mathcal{C}\}\setminus\{\{\}\}$
The property that alternating spacetime games have unique causal bridges (Dupuy’s terminology for enabling relations) translates to the same for the associated causal contextuality scenarios.
Definition 9 (Unique causal bridges).
Let $(X,O,\vdash,\mathcal{C})$ be a causal contextuality scenario. We say that the causal contextuality scenario has unique causal bridges if every measurement is on the right-hand-side of exactly one enabling relation (possibly with an empty left-hand side).
Lemma 1.
The causal contextuality scenario associated with an alternating spacetime game always has unique causal bridges.
Proof.
This is true by construction, since only one enabling relation is defined for every $x\in X$. ∎
Note that it is always possible to turn a causal contextuality scenario into a causal contextuality scenario that has unique causal bridges by duplicating and renaming measurements to make the right-hand sides of $\vdash$ unique, although it is not neutral in terms of semantics. This amounts to saying that, when a measurement experiment is performed (also with multiple contexts possible), the set of events that enabled this measurement is known to the experimenter, which is reasonable to assume in a local lab. However, a discussion of extending the isomorphism to non-unique causal bridges is given in Section 3. We focus our attention specifically to unique causal bridges in this paper.
Definition 10 (Game associated with a causal contextuality scenario).
Given a canonical causal contextuality scenario $(X,O,\vdash,\mathcal{C})$, we define the associated spacetime game with imperfect information as follows:
  • $\mathcal{P}=\{\text{Alfred},\text{Bob}\}$
  • Nodes played by Bob: $\mathcal{N}_{B}=\{t|\exists x,t\vdash x\}$
  • Nodes played by Alfred: $\mathcal{N}_{A}=\{(x,c)|x\in X,t\vdash x,c\in\mathcal{C}\cap\textsf{enabled}(t)\}$
  • $\mathcal{N}=\mathcal{N}_{A}\cup\mathcal{N}_{B}$
  • $\forall(x,c)\in\mathcal{N}_{A},\rho((x,c))=\text{Alfred}$
  • $\forall t\in\mathcal{N}_{B},\rho(t)=\text{Bob}$
  • Actions by Bob: $\mathcal{A}_{B}=\{\mathcal{C}\cap\textsf{enabled}(t)|t\in\mathcal{N}_{B}\}$
  • Actions by Alfred: $\mathcal{A}_{A}=\bigcup_{x}O_{x}$
  • $\mathcal{A}=\mathcal{A}_{A}\cup\mathcal{A}_{B}$
  • $\forall(x,c)\in\mathcal{N}_{A},\chi(x,c)=O_{x}$
  • $\forall t\in\mathcal{N}_{B},\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$
  • $\mathcal{R}_{A}=\{(t,(x,c))|t\vdash x\}$
  • $\mathcal{R}_{B}=\{((x,c),t)|x\in\textsf{domain}(t)\}$
  • $\mathcal{R}=\mathcal{R}_{A}\cup\mathcal{R}_{B}$
  • $\forall(t,(x,c))\in\mathcal{R}_{A},\sigma(t,(x,c))=c$
  • $\forall((x,c),t)\in\mathcal{R}_{B},\sigma((x,c),t))=t(x)$
  • Alfred’s information sets: $\mathcal{I}_{A}=\{(x,c)|x\in X\}\equiv X$
  • Bob’s information sets: $\mathcal{I}_{B}=\{\{t\}|t\in\mathcal{N}_{B}\}\equiv\mathcal{N}_{B}$
  • $\mathcal{I}=\mathcal{I}_{A}\cup\mathcal{I}_{B}$
Lemma 2.
The spacetime game associated with a canonical causal contextuality scenario with unique causal bridges is always alternating.
Proof.
The first property is true by construction.
The second property, that the graph is bipartite, holds by construction, as edges always connect a node in $\mathcal{N}_{A}$ to a node in $\mathcal{N}_{B}$ or vice versa.
The third property, that Bob’s information sets are singletons, also holds by construction since $\mathcal{I}_{B}=\mathcal{N}_{B}$.
The fourth property is that the nodes in the same information set of Alfred have a unique parent. These nodes are necessarily of the form $(x,c_{1})$, $(x,c_{2})$ for some measurement setting $x$ and contexts $c_{1}$ and $c_{2}$. By construction, the parent node of both of them must be some $t$ such that $t\vdash x$, and thus must be the same because we assume that the causal contextuality scenario has unique causal bridges.
The fifth property is that Bob’s information sets always have one successor. This is true because every one of them corresponds to the left hand side of one or more enabling relations, and the right-hand side of any of them gives a successor.
The sixth property is that all outgoing edges of the nodes $((x,c))_{c}$ in one of Alfred’s information sets $x$ have the same outgoing edges. This is true from the definition $\mathcal{R}_{B}=\{((x,c),t)|x\in\textsf{domain}(t)\}$ as well as of $\forall((x,c),t)\in\mathcal{R}_{B},\sigma((x,c),t))=t(x)$. Indeed, there is no dependency of any edge and label on the context.
The seventh property is that two distinct nodes played by Bob cannot have the same causal bridges. If we take two of these nodes, they correspond to two left-hand-sides of enabling relations $t$ and $s$. Let us further assume that the nodes for $t$ and $s$ have the same predecessors. This means by construction that $t$ and $s$ have the same domains, since the predecessors are made of the domain of $t$ plus contexts. Furthermore, if we assume that the labels of the edges from the predecessors are the same for $t$ and $s$, it implies that $t$ and $s$ exactly match on their full domain, also by construction of the labels, which are of the form $t(x)$ and $s(x)$ for each predecessor $x$. Thus, $t=s$ and the fourth property is fulfilled.
For the eigth property, let us take a node played by Bob (the left hand-side $t$ of some enabling relation), and that is connected to two nodes played by Alfred. These must be of the form $(x,c_{1})$ and $(y,c_{2})$ with $t\vdash x$ and $t\vdash y$. Let us also assume that the labels are identical. These labels correspond to the contexts in which $x$ and $y$ are measured, meaning that $x$ and $y$ are performed in the same context $c=c_{1}=c_{2}$. But then, $x$ and $y$ must be different, otherwise the nodes would be the same $(x,c)=(y,c)$.
Thus, the spacetime game associated with a causal contextuality scenario always fulfills the alternating property. ∎
Theorem 1.
The mapping from alternating spacetime games to canonical causal contextuality scenarios is injective.
Proof.
We prove that if a spacetime game $\mathcal{G}$ is mapped to a given canonical causal contextuality scenario $(X,O,\vdash,\mathcal{C})$, then $\mathcal{G}$ is necessarily (up to an isomorphism) the game associated with the causal contextuality scenario.
  • The players
    $\mathcal{P}=\{\text{Alfred},\text{Bob}\}$: this follows from the assumption that the game is alternating (first condition).
  • Alfred’s information sets
    It must be that $\mathcal{I}_{A}=X$ for the original game by construction of the causal contextuality scenario.
  • Bob’s nodes
    The left hand sides of the enabling relation are, by definition of the mapping from $\mathcal{G}$, exactly those of the form
    $t=\{(\iota(M),\sigma(M,P))|M\smile P\}$
    for each $P$ that is a node played by Bob in $\mathcal{G}$ and that is the parent of $x$ such that $t\vdash x$.
    Because of the alternating property, for any $t$ the corresponding node $P$ must be unique, as two of Bob’s nodes cannot have the same causal bridge (seventh condition). Conversely, any of Bob’s nodes must have a successor information set $x$ (also because of the alternating property, fifth condition) and thus, any of Bob’s nodes in the original game yields the left-hand side of an enabling relation. Thus, the correspondence between Bob’s nodes and the left-hand-sides of the enabling relations is exact: $\mathcal{N}_{B}=\{t|\exists x,t\vdash x\}$.
    This yields the pattern shown on Figure 7. It also follows that $\textsf{enabled}(t)=\textsf{successor}(t)$.
    Figure 7: The graph pattern implied by the definition of $t\vdash x$ in the mapping from games to causal contextuality scenarios. We show the pattern for two events in $t=\{(y,t(y)),(z,t(z))\}$ and two contexts for $x$ for $y$ and for $z$, but this extends to any number of events and any domain size.
  • Bob’s information sets
    $\mathcal{I}_{B}=\mathcal{N}_{B}$ necessarily because the original game is alternating (third condition).
  • Bob-Alfred edges
    Whenever $t\vdash x$ with $t$ being one of Bob’s nodes and $x$ one of Alfred’s information sets, the parent of all the nodes in $x$ must be $t$, because it exists and is unique over the entire information set according to the alternating property (fourth condition). It follows that edges from Bob to Alfred are of the form $(t,(x,c))$. Thus, $\mathcal{R}_{A}=\{(t,(x,c))|t\vdash x\}$ where the allowed values for $c$, which index the nodes in information set $x$, $((x,c))_{c}$, must still be defined.
  • Bob-Alfred labels
    $\forall(t,(x,c))\in\mathcal{R}_{A},\sigma(t,(x,c))=c$ follows from the renaming convention of the labels from Bob to Alfred. According to this convention, for any node $t$ played by Bob, there are as many outgoing edges from $t$ as there are pairs $(x,d)$ such that $x\in d$, $d\in\chi(t)$, and $\sigma(t,(x,d))=d$. Thus $d$ also indexes the nodes in the information set $x$ and can be identified with $c$, and the edge going from $t$ to node $(x,c)$ must be labeled with $\sigma(t,(x,c))=c$.
  • Bob’s actions by node
    We now show that $\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$. This amounts to showing that $\chi(t)=\mathcal{C}\cap\textsf{successor}(t)$
    Because of the naming convention of Bob’s action, we know that the labels in $\chi(t)$ are sets of information sets that are successors of $t$.
    Recall that $\mathcal{C}$ is by construction of the causal contextuality scenario a natural cover. Let us take an element $c$ in $\mathcal{C}\cap\textsf{successor}(t)$. By definition, $c$ is not the empty set. There is a corresponding complete history $z$ of the original game such that $\forall i\in\textsf{successor}(t),i\in c\iff z(i)\neq\bot$. This means that $c$ contains exactly the information sets that are activated by $t$ in $z$, and it there is at least one.
    Because of the naming convention of Bob’s nodes, it means that $c$ corresponds to the label born by the edges activating these information sets, and thus $c\in\chi(t)$.
    The converse is also true. If we take an action $c\in\chi(t)$, it labels the edges to all the nodes $(x,c)$ such that $x\in c$ because of the naming convention, and it is not empty because at least one successor exists. Thus, there exists a complete history assigning $c$ to $t$ (because the game is clean as has no unused nodes) and assigning each $x\in c$ to some value, and no value for the other successors of $t$. The active information sets of Alfred in this complete histories form a set that is an element $\hat{c}$ of $\mathcal{C}$, the natural cover, which is a superset of $c$. Thus $c=\hat{c}\cap\textsf{successor}(t)$, and $c$ is not the empty set. Thus, $c\in\mathcal{C}\cap\textsf{successor}(t)$.
    We have thus proven that $\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$.
    We also note that there cannot be any duplicates, i.e., there cannot be distinct nodes in $x$ indexed by the same $c\in\chi(t)$, because of the eight condition of the alternating property.
  • Alfred’s nodes
    We have shown that the nodes in an information set $x$ are indexed by $\chi(t)$ where $t\vdash x$, so that $\mathcal{N}_{A}=\{(x,c)|x\in X,t\vdash x,c\in\chi(t)\}$. Also, we have shown that $\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$. Thus $\mathcal{N}_{A}=\{(x,c)|x\in X,t\vdash x,c\in\mathcal{C}\cap\textsf{enabled}(t)\}$.
  • Alfred-Bob edges
    Since we have established the pattern on Figure 7, it follows that edges from Alfred to Bob connect each information set in the domain of $t$ to $t$. Thus, the edges from Alfred to Bob are given by $\mathcal{R}_{B}=\{((x,c),t)|x\in\textsf{domain}(t)\}$. All these edges must be present because of the alternating property’s (sixth condition) implying that the nodes in an information must have the same successors.
  • Alfred-Bob labels
    Again since we have established the pattern on Figure 7, the label on an from Alfred’s information set $y$ to Bob’s node $t$ must be $t(y)$. It follows that $\forall((x,c),t)\in\mathcal{R}_{B},\sigma((x,c),t)=t(x)$
  • Bob’s actions
    $\mathcal{A}_{B}=\{\mathcal{C}\cap\textsf{enabled}(s)|s\in\mathcal{N}_{B}\}$ because the game is clean and we already showed that $\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$.
  • Alfred’s actions by node
    $\forall(x,c)\in\mathcal{N}_{A},\chi(x,c)=O_{x}$. This is true by construction of the causal contextuality scenario.
  • Alfred’s actions
    : $\mathcal{A}_{A}=\bigcup_{x}O_{x}$. This is true by construction of the causal contextuality scenario.
  • Assignment to Alfred
    We have $\forall n\in\mathcal{N}_{A},\rho(n)=\text{Alfred}$ by construction of $\mathcal{N}_{A}$.
  • Assignment to Bob
    We have $\forall n\in\mathcal{N}_{B},\rho(n)=\text{Bob}$ by construction of $\mathcal{N}_{B}$.
  • All nodes
    We must have $\mathcal{N}=\mathcal{N}_{A}\cup\mathcal{N}_{B}$ because there are only two players according to the alternating property (first condition).
  • All edges
    $\mathcal{R}=\mathcal{R}_{A}\cup\mathcal{R}_{B}$ because there are only two players according to the alternating property (first condition).
  • All actions
    $\mathcal{A}=\mathcal{A}_{A}\cup\mathcal{A}_{B}$ because there are only two players according to the alternating property (first condition).
  • All information sets
    $\mathcal{I}=\mathcal{I}_{A}\cup\mathcal{I}_{B}$ because there are only two players according to the alternating property (first condition).
Theorem 2.
The mapping becomes bijective if we only consider causal contextuality scenarios with the natural cover of the associated alternating spacetime game.
Proof.
We observe that, in the other direction, if we have a causal contextuality scenario mapped to a spacetime game, then its components $X$ and $O$ must be given by our mapping. The $\vdash$ relation follows necessarily from the definition of the edges and labels. Only the cover may differ, however, in the special case that the cover is the natural cover of the corresponding spacetime, it is also recovered as well. Thus, we have an isomorphism from alternating spacetime games to causal contextuality scenarios with unique causal bridges, and back. ∎
2.6 The strategy sheaf of alternating games
Let us first consider the particular case of spacetime games with perfect information and the corresponding causal contextuality scenarios
If we take over the definition of a complete strategy of nature as a pure strategy appearing in the strategic form of the game (which is, indirectly, the strategic form of the associated game in extensive form with imperfect information) (Fourny, 2020), instead of that given in (Abramsky et al., 2024), we note that these strategies are simply partial functions on a discrete space ($\mathcal{I}$) to actions. Thus, we generally consider the restrictions of these partial functions to smaller domains, from which it follows that the suchly defined presheaf of strategies has the sheaf property. This was left as an open question by Abramsky et al. (2024) to determine “good covers” for which this is the case, and we have thus shown that the natural cover of the associated game in this paper is such a good cover.
The argument why the sheaf property obtains is identical to the one given by Abramsky and Brandenburger (2011) for the event sheaf, because it also involves the glueing of partial functions on discrete space that agree on their overlaps.
This result can naturally be extended to spacetime games with imperfect information: the conversion to the equivalent extensive form is made as for perfect information, and then, for each non-singleton information set in the spacetime game, the information sets corresponding in the extensive form to the individual nodes of the information set are merged.
Then, the pure strategies of nature for the game in extensive form is taken as a basis for the strategy sheaf.
2.7 Mixed strategies, Bell inequalities, and the strategy distribution presheaf
Nash equilibria (Nash, 1951) and other concepts apply naturally to spacetime games (Fourny, 2020). In particular, a mixed strategy in the Nash sense directly corresponds to a deterministic hidden variable model where $\lambda$ is a pure strategy. Free choice, as commonly assumed in the quantum foundations literature, directly translates to nature and the experimenters independently picking their mixed strategies – as is done when considering Nash equilibria in classical game theory.
In fact, if we observe that X and Y have disjoint supports (because Alice either picks basis x or y) and we merge them into a single one (call it X) – and we do the same for Bob (W and Z are merged into a single variable Z) – then the formula for computing the probability for the complete history resulting from a combination of a mixed strategy of nature (a distribution over the values of $\Lambda$) and the mixed strategies of experimenters (distributions over the values of $A$ and $B$) happens to exactly correspond to the formula used in the Bell inequality literature (Colbeck, 2017) for local-realist theories with measurement independence:
$P_{XZ|AB}=\sum_{\lambda}P_{\Lambda}(\lambda)P_{X|A\Lambda}P_{Z|B\Lambda}$
Figure 8 shows how this can also be interpreted as a causal network, with all variables depending deterministically on the mixed strategies, leading to a distribution of the complete histories for each combination of mixed strategies of Alice, Bob, and Alfred. Since in our non-Nashian paradigm (Fourny, 2019a)(Baczyk and Fourny, 2023) we posit a symmetry between nature and experimenters (or equivalently, between settings and outcomes), we add explicit lambdas also for the experimenters and not only for nature.
Figure 8: A visualization of the Nash mixed strategies as exogeneous nodes in a causal network. We merged XY and WZ into single random variables, by observing that the supports are disjoint, to keep the visual simple to understand.
In the special case of flat games (no adaptive measurements), it was shown by Abramsky and Brandenburger (2011) that the existence of a deterministic hidden variable model realizing an empirical model is implied by the existence of a global section for it. We make the observation that a mixed strategy of nature is the same mathematical object as a global section of the event distribution presheaf and we remember that spacetime games have a natural cover given by the projection of the support of their complete histories to nature. Then, one must conclude that the existence of a mixed strategy of nature leading to a distribution on the complete histories that is consistent with an empirical model (on the natural cover) implies the existence of a deterministic hidden variable model for this empirical model. This is consistent with our qualitative argument given by Baczyk and Fourny (2023) that Nashian game theory is not compatible with quantum physics.
3 Discussion
3.1 Non-unique causal pasts
It is straightforward to define spacetime games that correspond to causal contextuality scenarios with non-unique causal pasts: this is done by allowing for non-singleton information sets for Bob, with one node for each different causal past under which Bob carries out some experiments in his lab. However, we leave the extension of the isomorphism and the characterization of the corresponding subclasses for future work, as we expect that the characterization of these subclasses may be more complex.
3.2 Perfect information
Fourny (2020) observed in that imperfect information in the extensive form of the game corresponding to a spacetime game with perfect information can be interpreted in terms of the preservation of locality: information sets are grouped in the way they are because of the finite speed of light, and spacelike separation.
If we do not require perfect information for spacetime games, this introduces another physical justification for imperfect information in the resulting extensive form, namely, noncontextuality, because non-singleton information sets in the spacetime game (which lead to the fusion of the corresponding information sets in the extensive form) arise from the different contexts in which a measurement can be carried out, with the result independent of the context. This is also consistent with the fact that nonlocality and contextuality are related.
Having said this, we would like to conclude sharing our scientific belief, that only spacetime games with perfect information correspond to a physical reality, and that spacetime games with imperfect information are, above all, a useful mathematical tool for reductio ad absurdum proofs that nature is contextual. This is aligned with Dzhafarov et al. (2016) and on keeping locality as a fundamental principle of nature, provided we weaken measurement independence and assume non-Nashian free choice instead.
4 Related work
Hardy models (Hardy, 1993) fulfill the unique causal bridge constraints and are thus representable with spacetime games.
PR boxes (Popescu and Rohrlich, 1994) fulfill the unique causal bridge constraints and are thus representable with spacetime games.
GHZ models (Greenberger et al., 1989) fulfill the unique causal bridge constraints and are thus representable with spacetime games.
The framework also maps to contextuality by default (Dzhafarov et al., 2016), which corresponds to having perfect information. Forcing random variables corresponding to measurement outcomes to have the same value formally corresponds to grouping the corresponding nodes into the same information set.
Sheaves of sections as constructed by (Gogioso and Pinzani, 2021) correspond to a special case of spacetime games with perfect information in which measurements are performed independently of previous measurement outcomes, and without merging the information sets (see also Section 6.b. in (Abramsky et al., 2024)), which is also covered by the unique causal bridge condition. The fact that the sheaf property obtains is consistent with our results, which generalize the case covered in (Gogioso and Pinzani, 2021) to all alternating spacetime games, and causal contextuality scenarios with unique causal bridges.
A definition of free choice is given by Renner and Colbeck (2011), as a decision (called spacetime variable) being independent of anything it could not have caused. A deterministic resolution algorithm for all spacetime games with perfect or imperfect information, and which circumvents impossibility theorems (such as the sheaf property of the strategy distribution presheaf not generally holding) by weakening free choice, is given in (Fourny, 2019a), with the special case of strategic games covered in (Fourny, 2017) and for games in extensive form with perfect information in (Fourny et al., 2014). A high-level discussion of the non-Nashian paradigm in the context of quantum physics interpreted as a game between the experimenter and nature is given in (Fourny, 2019b).
The seeding philosophy of the non-Nashian approach can be found (Dupuy, 1992) (counterfactual decision theory) and the game theoretic seed and initial conjectures were made in (Dupuy, 2000).
There is also subsequent work by Kadnikov and Wichardt (2022) (also see (Kadnikov, 2020)) that provides a strategic decision making model generalizing the normal form and extensive form with graph structures related to spacetime games. This work defines the outcomes of the model based on sets of leaves of the graph structure and nicely connects it to the set-theoretic form by Bonanno (1992) (see also (Bonanno, 1991)). These sets of leaves can be put in correspondence with the complete histories, and the outcomes of the game in extensive form with imperfect information.
Hance and Hossenfelder (2022) give an excellent argument of why impossibility theorems do not disprove the existence of local theories. Hance et al. (2022) also is of the opinion that free choice can be weakened. Wharton and Argaman (2020) refers to this possibility as (non-)future-input-dependent parameters, and it is also known as measurement independence (see for example (Storz et al., 2023) for the explicit mention of this assumption in Bell experiments).
5 Acknowledgements
I thank my direct co-authors Jean-Pierre Dupuy, Stéphane Reiche, Michael Baczyk, Felipe Sulser, as well as my students at ETH who contributed to the game framework in various projects over the years: Emilien Pilloud, Ramon Gomm, Luc Stoffer, Florian Pauschitz. I also thank Michal Friedman, Richard Shiffrin, Renato Renner, V Vilasini, Philipp Kammerlander, Alexis Flesh, Lukas Sigrist, Philippe Colo, Abishek Ramdas, Günther Dissertori, Antoine Bommier, Gustavo Alonso, Daniel Kadnikov, Joseph Halpern, Robert French, Alexei Grinbaum, Matthew Jones, and VIS, for enlightening and stimulating discussions over the years and for their support.