### v3

# Spacetime games subsume causal contextuality scenarios

Abstract
We show that causal contextuality scenarios are equivalent to a subset of the formerly published spacetime games, which generalize game theory to decisions arbitrarily located in Minkowski spacetime. This insight leads to certain constructs and proofs being shorter, simpler, and more intuitive when expressed in the spacetime game framework than in the causal contextuality scenario framework. This shows that the insights of both frameworks taken together can contribute positively to advancing the field of quantum foundations.
keywords:
Game theory, Quantum theory, Contextuality, Locality, Bell inequalities, Special relativity
1 Introduction
Quantum theory has been challenging the physics community for a century now. The question remains open of whether there exists an extension theory, as envisioned by Einstein, Podolsky, and Rosen (EPR) [1], which is able not only to predict the same statistical distributions as the Born rule for repeated experiments—they are observed in Nature with very high precision—but also to predict each individual measurement outcomes.
But the way such a theory can look like is constrained. Indeed, two assumptions [2] are made for Bell inequalities to hold [^1] Determinism is derived from these two with a short argument made by Bell [2], thus the companion adjective “local-realist.”:
  • Einsteinian locality: no element of information propagates faster than light.
  • Measurement independence: the choice of a measurement setting, by the experimenter, is exogeneous and, as such, independent of anything not in its future light cone.
Since there is experimental evidence [3] that Bell inequalities are broken by Nature, it follows that one of these two assumptions does not hold. Most of the quantum physics literature so far has focused on assuming measurement independence and dropping Einsteinian locality. Little, but not nothing, has been published on the second possibility, namely, that measurement independence does not hold.
Perhaps this is related to the fact that this assumption of measurement independence is also commonly called “free choice” or “free will” and can lead to strong or emotional philosophical opinions on the matter: we are scientists, but we are also human beings. Yet, it is so that measurement independence is an assumption that can be formalized mathematically in terms of the statistical independence and/or the counterfactual independence (the two are distinct but related in subtle ways [4]) of random variables, and the scientific approach dictates that we must consider the possibility that measurement independence does not hold with as much reason as the possibility that it does. The risk of failing to do so is that an extension theory exists but remains undiscovered because we do not commit any resources to finding it.
One of the most interesting developments in the quantum foundations is the realization that many surprising and disturbing aspects of quantum theory can be abstracted away from physical experiments, and defined purely in terms of which experiments a physicist decides to carry out, and which outcomes they observe. Furthermore, one does not need infinite lists of choices: EPR setups are already interesting with two agents, two settings per agent, and two outcomes per agent and setting (2-2-2). We demonstrated [5] that one can model and view an experimental setup as a game between observers and the universe, Nature being an economic agent that minimizes action (principle of least action). Baczyk and Fourny [6] have given a qualitative argument that the Nash equilibrium [7] paradigm, which has been the mainstream framework of game theory for 75 years, corresponds to the conditions that lead to Bell inequalities:
  • locality corresponds to imperfect information (dashed lines);
  • realism corresponds to the strategic form that can be built from a game in extensive form;
  • measurement independence corresponds to unilateral deviations;
  • a local hidden-variable model corresponds to a mixed strategy (which is a statistical mixture of pure strategies) in the strategic form.
With this insight, one learns that any extension theory of quantum physics, with this game interpretation, cannot follow the Nash paradigm, and must follow instead a non-Nashian game theory paradigm, such as the Perfect Prediction Equilibrium or the Perfectly Transparent Equilibrium [4].
Subsequent work by Abramsky et al [8] builds on our idea that quantum experimental setups can be seen as games between observers and the universe and models quantum experiments as causal contextuality scenarios. In this paper, we show that our framework subsumes causal contextuality scenarios, i.e., that causal contextuality scenarios are equivalent to a subclass of spacetime games. We provide the formal mapping in both directions. We conclude by showing that the bridge between the two formalisms is mutually beneficial and leads to a clearer understanding of a decision-theoretical view of quantum physics. In particular, we discuss how certain seemingly complex constructs, reasonings, or proofs involving causal contextuality scenarios become one-line arguments in the spacetime game framework.
2 Background
But before we do so, we are first going through the spacetime game framework with a few examples.
2.1 Spacetime games and their extensive form
Game theory has two main game structures [9]: the normal form and the extensive form. As explained by Baczyk and Fourny [10][6], the normal form corresponds to spacelike-separated decisions, while the extensive form corresponds to timelike-separated decisions and even adaptive measurements. A general framework combining both, i.e., for decisions with arbitrary positions in spacetime, did not exist before 2019 to the best of our knowledge. If we are to consider generic quantum experimental setups with labs running local experiments placed at arbitrary positions in spacetime, and sending quantum states to one another—this corresponds to a fixed causal order in the process matrix framework [11]—, we need an extension and unification of the normal form and extensive form.
Thus, a new class of games called spacetime games with perfect information was introduced [5] [10]. It extends the semantics of game theory to special relativity. In spacetime games, the agents make decisions at various positions in Minkowski spacetime. Spacetime games can be seen as the least common denominator of games in normal form on the one hand (spacelike separation), and games in extensive form with perfect information on the other hand (timelike separation).
Figure 1: A spacetime game with perfect information that represents a Bell experiment with two spacelike-separated observers. Alice and Bob each select a choice of measurement context (here, only singleton contexts). Alice can pick X or Y, and Bob can pick W or Z. Then, for each one of the two measurements selected, Alfred selects an outcome (0 or 1). Perfect information in spacetime games is suitable for the study of non-locality, a special case of contextuality.
The same documents explain how these games can be used to model quantum experiments as a game played between observers and nature. An example of a game corresponding to the Bell experiment is shown in Figure 1, where Alice and Bob are the observers, and Alfred, whom we like to visualize as a cat, represents nature.
A canonical injection of the class of spacetime games with perfect information into the class of games in extensive form with imperfect information is also given. Figure 2 shows the game in extensive form with imperfect information that corresponds to the game of Figure 1. The two have equivalent semantics.
This is important because it means that spacetime games, when converted to games in extensive form with imperfect information, inherit several decades’ worth of research, constructs, and proofs carried out on games in extensive form with imperfect information for the past 75 years. This is key to our claim that several constructs and proofs behind causal contextuality scenarios can be significantly simplified. In other words, spacetime games provide an intuitive visual representation for quantum experimental setups but also a conversion algorithm to represent such setups as games in extensive form with imperfect information.
2.2 Imperfect information and the dashed lines
Imperfect information is visualized with dashed lines, which group nodes into equivalence classes called information sets. The meaning of these dashed lines is that the agent who makes the decision at these nodes does not know at which one of these nodes they are making their decision: they are not fully informed about the decisions made at ancestor nodes [^2] These ancestor nodes need not all be in the past, they can also be spacelike-separated in the spacetime game. Since information cannot propagate faster than light, it is not possible to be informed about decisions made at spacelike-separated locations. in the extensive form. This is why it is called imperfect information. Concretely, in Figure 2, Bob is not informed of Alice’s decision ($\{X\}$ or $\{Y\}$), and this is why his two nodes are connected with a dashed line. This has to do with relativity theory: although Alice may have already made her decision, the information has not reached Bob yet because his decision is spacelike-separated from Alice’s, as can be seen in the spacelike form of Figure 1, which is a causal dependency graph in the sense of relativity theory.
The information sets of the extensive form in Figure 2 correspond exactly to those of the spacetime game in Figure 1 (which are singletons, one for each node because the spacetime game has perfect information). This is always the case for any spacetime games and their extensive form. The generalization of this canonical injection to all spacetime games, even spacetime games with imperfect information, is straightforward and explained in this paper.
2.3 Imperfect information, locality, and non-contextuality
The semantics of imperfect information and dashed lines have a subtle difference depending on whether one looks at the spacetime game or its extensive form. In a spacetime game with imperfect information, dashed lines physically represent non-contextuality, connecting identical decisions made in different contexts, whereas in a game in extensive form with imperfect information (at least those built from a spacetime game), dashed lines represent either locality or non-contextuality. The spacetime game framework makes explicit that locality is a special case of non-contextuality: in local scenarios, the spacetime game has perfect information and the extensive form has imperfect information, while in more general non-contextuality scenarios, both of them have imperfect information.
Figure 2: The game in extensive form with imperfect information that corresponds to the spacetime shown in Figure 1. Alice and Bob have each one information set (A and B) for their choice of measurement setting. Alfred has four information sets X, Y, W, and Z, one for each measurement. For scenarios of interest to the quantum foundations, i.e., exhibiting locality or non-contextuality, the game in extensive form will typically have imperfect information, regardless of whether the spacetime game has perfect information (for locality) or imperfect information (for more general cases of non-contextuality).
Because of the canonical injection of spacetime games into the class of games in extensive form with imperfect information, many well-studied concepts in the game theory literature apply to spacetime games. This includes flattening the strategic space with the (reduced) strategic form of a spacetime game [10][12], computing Nash equilibria [7], mixed strategies (which we will see are also global sections of empirical models), etc.
2.4 Causal contextuality scenarios
Subsequent work by Abramsky et al [8] suggests modeling a subclass of such games, which they call causal contextuality scenarios, as strategy presheaves. As we show in this paper, when natural covers are taken, causal contextuality scenarios are isomorphic to a subclass of spacetime games, which we call alternating spacetime games. This means that spacetime games recover causal contextuality scenarios as a special case, and by transitivity (based on the claims made in [8]), that they also recover the framework of Gogioso and Pinzani as a special case [13], as well as Hardy models [14], PR boxes [15], and GHZ models [16].
We also show that a flattened definition of the strategy presheaf based on reduced strategic forms [12] emerges naturally and that the sheaf property holds, and that temporal contextuality scenarios are subsumed by games in extensive form with perfect information.
3 Spacetime games
3.1 Definition
To define the mapping, we now give a formal definition of spacetime games. It is equivalent to that previously published [5] [10], but in a more compact form (see also [6]). In this definition, we do not require perfect information. Imperfect information has the meaning proposed by Kuhn [12] and extends naturally to spacetime games without affecting the injection into the class of games in extensive form with imperfect information. [^3] This is because merging nodes into information sets in the spacetime game translates into merging the corresponding information sets in the extensive form. Spacetime games with perfect information allow for the study of locality, while spacetime games with imperfect information allow for the more general study of non-contextuality.
Definition 1 (Spacetime game).
A spacetime game $\mathcal{G}$ is a tuple $(\mathcal{N},\mathcal{R},\mathcal{P},\mathcal{A},\rho,\chi,$ $\sigma,\mathcal{I},\mathcal{Z},u)$ where
  • $(\mathcal{N},\mathcal{R})$ is a directed acyclic graph of decision nodes, $\mathcal{N}$ being the nodes and $\mathcal{R}$ being the edges.
  • $\mathcal{P}$ is a set of players and $\mathcal{A}$ is a set of actions (in our context, actions are measurement settings and measurement outcomes)
  • Nodes are labeled with players: $\rho\in\mathcal{P}^{\mathcal{N}}$
  • Nodes are labeled with sets of actions: $\chi\in\mathcal{P}(\mathcal{A})^{\mathcal{N}}$
  • Edges are labeled with actions: $\sigma\in\mathcal{A}^{\mathcal{R}}$ labels each edge $(N,M)$ with an action $\sigma(N,M)\in\chi(N)$.
  • A partition of $\mathcal{N}$ into information sets: a set $\mathcal{I}$ of subsets of $\mathcal{N}$ that form a partition, in a way that is compatible with $\rho$ and $\chi$, meaning that
    $\forall i\in\mathcal{I},\forall N,M\in i,\rho(N)=\rho(M)\wedge\chi(N)=\chi(M)$
  • $\mathcal{Z}\in(\mathcal{A}\cup\{\bot\})^{\mathcal{I}}$ is a set of outcomes, also called complete histories, which are partial functions from information sets to actions. They must obey some constraints implied by the game structure. These constraints are defined further down in Section 3.4. The domain of definition, also called path, of an outcome is the set of information sets not mapped to $\bot$, which is the symbol used for unassigned actions.
  • $u=(u_{i})_{i\in P}$ with, for each $i$, $u_{i}\in\mathbb{R}^{\mathcal{Z}}$ is a family of utility functions mapping outcomes to payoffs, with one payoff for each outcome and player.
The payoffs associated with the outcomes are part of a game’s semantics but are not the focus of the present paper, which focuses on the game structure. They are also not part of any causal contextuality scenario and are ignored in the bijective mapping.
3.2 Further examples
Figure 3 shows an example of a spacetime game with imperfect information (the dashed lines have the same meaning as in Figure 2) and Figure 4 shows the corresponding extensive form with imperfect information.
Figure 3: A spacetime game with imperfect information that represents a non-contextuality experiment with two settings to pick out of three. Bob has one information set with three actions (the three contexts XY, XZ, and YZ). Alfred has 3 information sets X, Y, and Z for each measurement, with two actions each, 0 and 1. This is also known in the literature as a cyclic system of rank 3. Imperfect information in spacetime games captures the more general case of non-contextuality than the special case of locality.
Note that, to ease visualization, actions on the leaf nodes are shown as “virtual” edges with no destination node and labeled with the possible actions at these leaf nodes, but these edges are purely visual and do not belong to $\mathcal{R}$; instead, the edges leaving a leaf node $N$ correspond to the elements of $\chi(N)$.
Figure 4: The game in extensive form with imperfect information that corresponds to the spacetime game shown in Figure 3. It has the same information sets: XY, YZ, and XZ for Bob, and X, Y, and Z for Alfred.
Figure 5 shows an example of a spacetime game with perfect information that has a causal bridge (adaptive measurement) and the corresponding extensive form is shown in Figure 6. Spacetime games can be much more complex and mix spacelike-separation as in Figure 1, timelike-separation as in Figure 5, and imperfect information for explicit contexts as shown in Figure 3.
Figure 5: A spacetime game with imperfect information that shows an example of causal bridge: Bob only carries out his experiment if Alice measured Y and obtained an outcome of 0.
Figure 6: The game in extensive form with, in this case, perfect information, that corresponds to the spacetime game shown in Figure 5. It happens to have the same tree shape because there is no spacelike separation. Such games correspond to temporal contextuality scenarios.
Since locality is a special case of non-contextuality, the Bell experiment can also be represented with a single player picking a context with two measurements, in a cyclic manner (XW, WY, YZ, ZX). The corresponding game is shown in Figure 7. The corresponding game in extensive form is shown in Figure 8 and it is almost the same as the game shown in Figure 2, with the observer players merged into one, which amounts to looking at the Cartesian product of the Hilbert spaces.
Figure 7: A spacetime game with imperfect information that represents a Bell experiment as a non-contextuality experiment instead of a locality experiment, with a single observer player picking a context (A) of two measurement settings out of four possible contexts, and four information sets (X, Y, W, Z) played by nature. This is also known in the literature as a cyclic system of rank 4.
Figure 8: The game in extensive form with imperfect information that corresponds to the spacetime game shown in Figure 7. It has the same information sets: A for Bob, and X, Y, W, and Z for Alfred.
3.3 Notations
In the remainder of this paper, whenever we refer to a game $\mathcal{G}$, we will also implicitly use the above notations ($\mathcal{N},\mathcal{R}$, $\mathcal{P}$, $\mathcal{A}$, $\rho$, $\chi$, $\sigma$, $\mathcal{I}$) without explicitly defining them every time. This is because we only ever consider one game at a time, and it eases the read.
We write $N\smile M$ as a convenient shorthand for $(N,M)\in\mathcal{R}$.
We define for any node N
$\textsf{precedessors}(N)=\{M\in\mathcal{N}|M\smile N\}$
$\textsf{successors}(N)=\{M\in\mathcal{N}|N\smile M\}$
We also define for any information set $i$
$\textsf{precedessors}(i)=\bigcup_{N\in i}\textsf{predecessors}(N)$
$\textsf{successors}(i)=\bigcup_{N\in i}\textsf{successors}(N)$
We also use a slight abuse of notation whenever all nodes in the same information set share a common property: for example, $\rho(i)$ and $\chi(i)$ where $i$ is an information set. We also use, for convenience, a function $\iota\in\mathcal{I}^{\mathcal{N}}$ that maps each node to the information set it belongs to.
Conversely, we will write $z(N)$ for $z(\iota(N))$ where $z$ is an outcome (complete history) and $N$ is a node.
When we write the statement $\sigma(N,M)=a$ for some $a$, it implicitly implies $N\smile M$, which allows more concise formulas.
3.4 Histories
We now give a compact definition of the histories of a spacetime game (which is a compact form of that given in [10][6]):
Definition 2 (Histories).
Given a spacetime game in extensive form $\mathcal{G}$, a history is an assignment from $\mathcal{I}$ to $\mathcal{A}\cup\{\bot\}$ that only assigns an action to activated information sets. An information set is activated if it has a node such that the label of each one of its incoming edges is equal to the action assigned (by that same history) to the information set containing the source node of that edge:
$\forall i\in\mathcal{I},(z(i)\neq\bot\implies\exists N\in i,\forall M\in\textsf{predecessors}(N),z(M)=\sigma(M,N))$
The set of all histories is denoted $\mathcal{H}$.
In [10], we use the term “contingency coordinates” of a node $N$ to denote the minimal history not containing $N$ that coincides with any history in which $N$ is assigned an action (i.e., not $\bot$). In this paper, we use the term “causal bridge” [17] [^4] The idea of a decision happening only if a certain causal past is secured, and otherwise not, can be traced back to Dupuy [18][19], and we retain his terminology “causal bridge”. of a node $N$ to specifically designate the predecessors of $N$ and the labels on the edges from the predecessors to $N$. The two differ in that the contingency coordinates involve the entire ancestry, while the causal bridge only involves (directly connected, one level up) predecessors.
A history is complete if the implication becomes an equivalence. In a valid game, the set of outcomes is the set of complete histories.
Definition 3 (Complete histories, and validity of a spacetime game).
A spacetime game in extensive form $\mathcal{G}$ is valid if the set of outcomes $\mathcal{Z}$ matches the set of complete histories. A complete history is defined as a history that assigns an action to all the activated information sets.
$\forall i\in\mathcal{I},(z(i)\neq\bot\iff\exists N\in i,\forall M\in\textsf{predecessors}(N),z(M)=\sigma(M,N))$
We always assume that spacetime games are valid, as we consider validity to be part of their definition. Given a complete history (outcome) $z$, we write:
$\textsf{domain}(z)=\{i\in\mathcal{I},z(i)\neq\bot\}$
Without loss of expressive power, we assume that games do not contain unused information sets, i.e., there is no information set that never gets activated; otherwise, unused information sets can be pruned from the game. [^5] The same assumption will be made for causal contextuality scenarios: There is no measurement that never gets activated, otherwise such measurements can be pruned from the causal contextuality scenario. Likewise, we assume there are no unused actions in $\mathcal{A}$.
4 Results
We now formally show that, for a subclass of two-player spacetime games (which we call “alternating”), we can build an equivalent causal contextuality scenario. We also show that, in the other direction, we can convert back causal contextuality scenarios to two-player spacetime games.
Thus, spacetime games and causal contextuality scenarios stand in bijection for a broad class of experimental protocols, which can be either characterized as alternating spacetime games or as causal contextuality scenarios with “good covers” (see open question 7.3 by Abramsky et al [8]), whereby we suggest that a cover is “good” is it corresponds to the natural cover implied by a spacetime game structure.
Complete histories under the two frameworks are also in bijection, and histories in causal contextuality scenarios correspond to those histories in spacetime games that have a domain with even cardinality (in other words, incomplete histories in causal contextuality scenarios always stop with nature giving an outcome for the chosen measurement, and never end abruptly with a measurement chosen but not carried out).
4.1 Alternating spacetime games
Spacetime games, but also games in normal or extensive form with or without perfect information, are very general and cover many different scenarios, from microeconomics to leisure (Chess, etc). Here, we are looking at specific kinds of spacetime games in which experimenters and the universe alternate in choosing settings and outcomes. We thus now give a characterization of a particular relevant subclass of spacetime games called “alternating spacetime games”, for which we will show that there is a mapping with the corresponding causal contextuality scenarios.
Note that, since causal contextuality scenarios do not specify which observer performs which experiment, we call all our observers Bob for the mapping. If a causal contextuality scenario is extended with information on the identity of the observers (e.g. Alice, Bob) then they can be marked accordingly in the corresponding game.
In the definition of the alternating property, we interpret the game’s assumed features in terms of quantum-experimental protocols, to facilitate the understanding of the mapping that will be given further down.
Definition 4 (Alternating properties).
A spacetime game $\mathcal{G}$ is an alternating game if it is valid and it fulfills the following alternating properties:
  1. 2-PLAYERS
    It has two players (we call them Alfred, who is nature, and Bob, who is the observer).
  2. BIPARTITE
    The graph $(\mathcal{N},\mathcal{R})$ always connects a node played by Alfred to a node played by Bob, or a node played by Bob to a node not played by Alfred. In other words, the graph structure of the spacetime game is a bipartite graph if we ignore the direction of the edges.
  3. EVEN
    All root nodes are played by Bob. All leaf nodes are played by Alfred.
  4. BOB-S
    All information sets played by Bob are singletons. It means Bob is fully informed about decisions in his causal past, even in situations in which he can carry out the same experiment under different circumstances.
  5. BOB-A
    At any of Bob’s nodes, all available actions are used on at least one edge. [^6] One can also consider allowing for at most one unused action in $\chi(N)$ when looking at causal contextuality scenarios with contexts in which the antichain property does not hold.
    $\forall N\in\rho^{-1}(\text{Bob}),\forall a\in\chi(N),\exists M\in\mathcal{N},\sigma(N,M)=a$
  6. BA1
    Each node played by Alfred has exactly one parent node:
    $\forall N\in\mathcal{N},\rho(N)=\text{Alfred}\implies\exists!P\in\mathcal{N},P\smile N$
  7. BA2
    Given a node $N$ played by Bob, two nodes (played by Alfred) connected to $N$ with the same label (it is a measurement context) must be in different information sets (these are measurement settings). In other words, distinct nodes for the same measurement in the same context would be superfluous.
    $\forall N\in\rho^{-1}(\text{Bob}),\forall S,T\in\textsf{successors}(N),\sigma(N,S)=\sigma(N,T)\implies\iota(S)\neq\iota(T)$
  8. AB1
    All nodes in the same information set played by Alfred [^7] This trivially applies to Bob, too. have the same outgoing edges: same labels, same destination nodes. In other words, the causal future of a measurement does not depend on the context in which it was carried out.
    $\forall i\in\rho^{-1}(\text{Alfred}),\forall N,M\in i,\textsf{successors}(N)=\textsf{successors}(M)$
    $\forall i\in\rho^{-1}(\text{Alfred}),\forall N,M\in i,\forall T\in\textsf{successors}(i),\sigma(M,T)=\sigma(N,T)$
  9. AB2
    Two distinct nodes played by Bob cannot have the same causal bridge.
    $\forall N,M\in\rho^{-1}(\text{Bob}),$
    $(\textsf{predecessors}(N)=\textsf{predecessors}(M)\wedge$
    $\forall P\in\textsf{predecessors}(N),\sigma(P,N)=\sigma(P,M))\implies N=M$
Remark 1.
The height of an alternating spacetime game is always even, nodes at even depths (0, 2, …) are played by Bob, and nodes at odd depths (1, 3, …) are played by Alfred.
4.2 Non-locality as a special case of contextuality
AB2 amounts to saying that we assume that the measurements carried out by Bob under the same enabling circumstances are grouped into a unique experiment where Bob has multiple contexts to choose from. Note that this does not cause any loss of generality because this amounts to saying that non-locality is a special case of contextuality. See for example Figure 7 (which satisfies AB2), which groups the local experiments of Figure 1 (which does not satisfy AB2) into a single one.
The purpose of alternating spacetime games is to characterize a subset of spacetime games that is isomorphic to causal contextuality scenarios with good covers. When using the spacetime game framework for discussing non-locality, it is more desirable to consider the (factored) game in Figure 1 rather than the (expanded) game in Figure 7, knowing that the factored game can be considered equivalent to the same causality contextuality scenario as the one associated with the expanded game by the mapping.
4.3 Standardization of Bob’s actions in alternating spacetime games
Without loss of generality, we can put constraints on the names given to Bob’s actions in an alternating spacetime game to facilitate their interpretation as contexts, like so:
  • Each edge from a node played by Bob to a node played by Alfred is labeled with a non-empty set (the context) of information sets (the settings) played by Alfred.
  • For any node $N$ played by Bob, there are as many outgoing edges from $N$ as there are context-setting pairs $(x,c)$ such that $x\in c$ and $c\in\chi(N)$. The edge corresponding to $(x,c)$ must go to a node $M\in x$ and must be labeled with $\sigma(N,M)=c$.
In other words, Bob’s actions are contexts, and for each context, there are edges to each measurement setting in that (possibly empty) context, labeled with that context.
It is always possible to rename Bob’s actions such that this is the case: Given a node $N$ played by Bob and an action $c\in\chi(N)$, we rename $c$ to $\{\iota(M)|M\in\mathcal{N}\wedge\sigma(N,M)=c\}$. Because of BA2, the renaming is bijective.
We show below a correspondence table between the two frameworks in Table 1.
Spacetime games Causal contextuality scenarios
Alfred The nature player
Bob The experimenter
An information set played by Alfred A measurement
A node played by Alfred A measurement and its context
An information set played by Bob A set of events, which is the left-hand-side of an enabling relation
An action by Alfred A measurement outcome
An action by Bob A measurement context
An edge from Bob to Alfred Bob picks a measurement context and carries out a measurement in this context.
The label is the context.
An edge from Alfred to Bob An outcome $o$ is obtained for a specific measurement $x$, and is part of enabling
conditions $t$ for another measurement. The label is $t(x)=o$
A complete history A complete history
A history with an even number of defined decisions A history
An alternating game A causal contextuality scenario
An alternating game that is in normal form A flat contextuality scenario
An alternating game whose extensive form has perfect information A temporal contextuality scenario
The reduced strategic form of the spacetime game Flattening construction
A pure strategy of the reduced strategic form A global strategy
A mixed strategy of the reduced strategic form The global section of an empirical model (if it exists)
(Dupuysian) causal bridge Causally secured
The projection of a complete history to Bob A “good’) cover
Table 1: The correspondence of terminology between alternating spacetime games and causal contextuality scenarios.
4.4 Causal contextuality scenarios
We recall the definition of a causal contextuality scenario as given by Abramsky et al [8]:
Definition 5 (Causal contextuality scenario).
A causal contextuality scenario $\Gamma$ is a quadruplet $(X,O,\vdash,\mathcal{C})$ where X is a set of measurement settings, $O=(O_{x})_{x\in X}$ is a family of sets of measurement outcomes, indexed by settings, $\vdash$ is an enabling relation between a section on a subset of $X$ (a “consistent set of events”) and settings, and $\mathcal{C}$ is a family of subsets of $X$ whose union is X.
In this paper, we consider “clean” scenarios in the sense that there are no unused measurement settings, i.e., settings that cannot be enabled under any circumstances under the considered cover. If a causal contextuality scenario contains an unused setting, then it can be removed without affecting its semantics. This is analogous to our assumption that spacetime games do not have unused information sets.
Furthermore, this paper restricts its attention to covers that satisfy the antichain property. It is, however, possible to extend the mapping naturally even if the antichain property does not hold (e.g., by allowing unused actions at Bob’s nodes).
4.5 Natural cover
Each spacetime game (not necessarily alternating) has a natural cover that corresponds to all combinations of measurements that are performed in one of the complete histories:
Definition 6 (Natural cover associated with a game).
Given a spacetime game $\mathcal{G}$ with nature player Alfred, the natural cover $\mathcal{C}$ is defined as follows: For each outcome $z\in\mathcal{Z}$, we take the subset $C_{z}=\{i\in\textsf{domain}(z)|\rho(i)=\text{Alfred}\}$ of its domain of definition with only and exactly the information sets that are played by Alfred. Then the natural cover is $\mathcal{C}=\{C_{z}|z\in\mathcal{Z}\}$. Note that duplicates are implicitly eliminated because of the definition of a set.
For example, the natural cover of the game shown in Figure 1 (or 7) and Figure 2 (or 8) is $\{\{XW\},\{XZ\},\{YW\},\{YZ\}\}$, which is the cycle of rank 4.
The natural cover of the spacetime game shown in Figure 3 and Figure 4 is $\{\{XY\},\{XZ\},\{YZ\}\}$, which is the cycle of rank 3.
The natural cover of the spacetime game shown in Figure 5 and Figure 6 is $\{\{X\},\{YW\},\{YZ\}\}$.
This natural cover will be used to define the corresponding measurement scenario for any alternating spacetime game.
4.6 The mapping
4.6.1 Causal contextuality scenario associated with an alternating spacetime game
Any alternating spacetime game can be associated with a causal contextuality scenario as follows. As explained in the previous tabular correspondence, the general idea of the overall mapping is that:
  • Alfred’s information sets are measurements;
  • Bob’s (singleton) information sets are the left-hand sides of the enabling relations of the causal contextuality scenario;
  • Bob’s actions (which label edges from Bob to Alfred) are contexts. Alfred’s nodes within the same information set are measurements with the same measurement setting but in different contexts and possibly in different causal pasts (i.e., different parent nodes);
  • Alfred’s actions (which label edges to Bob) are the measurement outcomes.
We now give a formal definition of the above:
Definition 7 (Causal contextuality scenario associated with an alternating spacetime game).
Given an alternating spacetime game $\mathcal{G}$ with nature player $\text{Alfred}\in\mathcal{P}$, we define the equivalent causal contextuality scenario $\Sigma(\mathcal{G})=(X,O,\vdash,\mathcal{C})$ as follows:
  • $X=\{i\in\mathcal{I}|\rho(i)=\text{Alfred}\}$
  • for each $x\in X$, $O_{x}=\chi(x)$
  • for each $x\in X$, and for each $P\in\textsf{predecessors}(x)$, we define that $t_{P}\vdash x$ where
    $t_{P}=\{(\iota(M),\sigma(M,P))|M\smile P\}$
  • additionally, for each $x\in X$, if there is a node in $x$ with no parent, we also define {}$\{\}\vdash x$.
  • $\mathcal{C}$ is the natural cover of $\mathcal{G}$.
For it to be a valid causal contextuality scenario, we must prove that the enabling relations have consistent left-hand sides.
Lemma 1.
The causal contextuality scenario associated with an alternating spacetime game always has consistent sets of events on the left-hand sides of enabling relations.
Proof.
AB1 implies that two nodes in $y$, which is played by Alfred, have the same labels and destination nodes. As a consequence, if we have two nodes $M,N\in y$, it follows that $\sigma(M,P)=\sigma(N,P)$, leading to only one assignment for $t(y)$. ∎
Remark 2.
It follows from the lemma, given some $x\in X$ and one of its predecessors $P$, that $\sigma(M,P)=t_{P}(y)$ for any $M\in y$. See also Figure 9 for a visual of this pattern.
Figure 9: The graph pattern implied by the definition of $t\vdash x$ in the mapping from games to causal contextuality scenarios. We show the pattern for two events in $t=\{(y,t(y)),(z,t(z))\}$ two contexts for $y$ and $z$, and four contexts (among which two under causal bridge $t$) for $x$, but this extends to any number of events and any domain size.
Lemma 2.
$P\mapsto t_{P}$ is a bijection between the set of nodes played by Bob and the left-hand sides of the mapped enabling relations. More precisely, for any $x\in X$, the left-hand sides of the enabling relations to $x$ in the causal contextuality scenario are in bijection with the predecessors of $x$ in the alternating spacetime game.
Proof.
Every $P$ leads to at least one enabling relation with left-hand-side $t_{P}$ because it has at least one child (property EVEN). It follows from the property AB2 that $t$ is injective. Since the codomain is defined as the range, the mapping is also surjective. ∎
We use the abbreviation $\textsf{enabled}(t)=\{x\in X|t\vdash x\}$ and use this notation in particular for restricting the cover $\mathcal{C}$ to the “local cover at t” denoted $\mathcal{C}_{t}$ which we define as
$\mathcal{C}_{t}=\{C\cap\textsf{enabled}(t)|C\in\mathcal{C}\}\setminus\{\emptyset\}$
Note that $\mathcal{C}_{t}$ cannot be empty because there are no unused settings, and it does not contain the empty set.
4.6.2 Alternating spacetime game associated with a causal contextuality scenario
We now give the mapping in the opposite direction, i.e., how to associate a spacetime game with a causal contextuality scenario.
Definition 8 (Game associated with a causal contextuality scenario).
Given a causal contextuality scenario $(X,O,\vdash,\mathcal{C})$, we define the associated spacetime game with imperfect information obtained like so:
  • $\mathcal{P}=\{\text{Alfred},\text{Bob}\}$
  • Nodes played by Bob: $\mathcal{N}_{B}=\{t|\exists x\in X,t\vdash x\}$
  • Nodes played by Alfred: $\mathcal{N}_{A}=\{(t,x,c)|t\vdash x,c\in\mathcal{C}_{t}\}$
  • $\mathcal{N}=\mathcal{N}_{A}\cup\mathcal{N}_{B}$
  • $\forall(t,x,c)\in\mathcal{N}_{A},\rho((t,x,c))=\text{Alfred}$
  • $\forall t\in\mathcal{N}_{B},\rho(t)=\text{Bob}$
  • Actions by Bob: $\mathcal{A}_{B}=\bigcup_{t\in\mathcal{N}_{B}}\mathcal{C}_{t}$
  • Actions by Alfred: $\mathcal{A}_{A}=\bigcup_{x\in X}O_{x}$
  • $\mathcal{A}=\mathcal{A}_{A}\cup\mathcal{A}_{B}$
  • $\forall(t,x,c)\in\mathcal{N}_{A},\chi(t,x,c)=O_{x}$
  • $\forall t\in\mathcal{N}_{B},\chi(t)=\mathcal{C}_{t}$
  • $\mathcal{R}_{A}=\{(t,(t,x,c))|t\vdash x\}$
  • $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{domain}(t)\}$
  • $\mathcal{R}=\mathcal{R}_{A}\cup\mathcal{R}_{B}$
  • $\forall(t,(t,x,c))\in\mathcal{R}_{A},\sigma(t,(t,x,c))=c$
  • $\forall((u,x,c),t)\in\mathcal{R}_{B},\sigma((u,x,c),t))=t(x)$
  • Alfred’s information sets: $\mathcal{I}_{A}=\{(t,x,c)|x\in X\}\equiv X$
  • Bob’s information sets: $\mathcal{I}_{B}=\{\{t\}|t\in\mathcal{N}_{B}\}\equiv\mathcal{N}_{B}$
  • $\mathcal{I}=\mathcal{I}_{A}\cup\mathcal{I}_{B}$
  • $\mathcal{Z}$ is as defined in Definition 3 to make the game valid.
Lemma 3.
The spacetime game associated with a causal contextuality scenario is alternating.
Proof.
The game is valid by construction. Thus, we need to show that it is alternating.
2-PLAYERS is true by construction since $\rho$ can only take two values.
BIPARTITE holds by construction, as edges always connect a node in $\mathcal{N}_{A}$ to a node in $\mathcal{N}_{B}$ or vice versa.
EVEN holds by construction: any node $(t,x,c)$ played by Alfred has $t$ as its parent node, so all root notes must be played by Bob. Likewise, given a node $t$ played by Bob, there is at least one enabling relation $t\vdash x$ that it was taken from. Furthermore, $\mathcal{C}_{t}$ is never empty, so as a consequence, $t$ has at least one child node.
BOB-S (Bob’s information sets are singletons) also holds by construction since $\mathcal{I}_{B}=\mathcal{N}_{B}$.
BOB-A (no unused action at Bob’s nodes) is true: given a node $t$ played by Bob and an action at this node, i.e., a context $c$ in (non-empty) $\mathcal{C}_{t}$, $c$ is used on the edge from $t$ to $x$ for any $x\in c$.
BA1 (nodes played by Alfred have exactly one parent) is true by the construction of $\mathcal{R}_{A}$ because the parent of $(t,x,c)$ is necessarily $t$.
For BA2, let us take a node played by Bob (the left-hand side $t$ of some enabling relation), and that is connected to two nodes played by Alfred. These must be of the form $(t,x,c_{1})$ and $(t,y,c_{2})$ with $t\vdash x$ and $t\vdash y$. Let us also assume that the labels are identical. These labels correspond to the contexts in which $x$ and $y$ are measured, meaning that $x$ and $y$ are performed in the same context $c=c_{1}=c_{2}$. But then, $x$ and $y$ must be different, otherwise the nodes would be the same $(t,x,c)=(t,y,c)$.
AB1 requires that all outgoing edges of the nodes $((u,x,c))_{u,c}$ in one of Alfred’s information sets $x$ have the same outgoing edges. This is true from the definition $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{domain}(t)\}$ as well as of $\forall((u,x,c),t)\in\mathcal{R}_{B},\sigma((u,x,c),t))=t(x)$. Indeed, there is no dependency of any edge and label on the context.
AB2 requires that two distinct nodes played by Bob cannot have the same causal bridges. If we take two of these nodes, they correspond to two left-hand sides of enabling relations $t$ and $s$. Let us further assume that the nodes for $t$ and $s$ have the same predecessors. This means by construction that $t$ and $s$ have the same domains since the predecessors are made of the domain of $t$ plus contexts. Furthermore, if we assume that the labels of the edges from the predecessors are the same for $t$ and $s$, it implies that $t$ and $s$ exactly match on their full domain, also by construction of the labels, which are of the form $t(x)$ and $s(x)$ for each predecessor $x$. Thus, $t=s$ and BA1 is fulfilled.
Thus, the spacetime game associated with a causal contextuality scenario is always alternating. ∎
Theorem 4.
The mapping from alternating spacetime games to causal contextuality scenarios as defined in Definition 7 is injective.
Proof.
We prove that if a spacetime game $\mathcal{G}$ is mapped to a given causal contextuality scenario $(X,O,\vdash,\mathcal{C})$ according to Definition 7, then $\mathcal{G}$ is necessarily, up to an isomorphism [^8] We mean here up to renaming the player names, the labels, etc without affecting the semantics of the game., the game associated with the causal contextuality scenario according to Definition 8.
  • The players
    $\mathcal{P}=\{\text{Alfred},\text{Bob}\}$: this follows from property 2-PLAYERS.
  • Alfred’s information sets
    It must be that $\mathcal{I}_{A}=X$ for the original game by construction of the causal contextuality scenario (first bullet point of Definition 7).
  • Bob’s nodes
    The left-hand sides of the enabling relation are, by definition of the mapping from $\mathcal{G}$, exactly those of the form
    $t(P)=\{(\iota(M),\sigma(M,P))|M\smile P\}$
    for each $P$ that is a node played by Bob in $\mathcal{G}$ and that is the parent of some $M\in x$ such that $t\vdash x$.
    Lemma 2 showed that this is a bijection (because of AB2 and EVEN) and that $\mathcal{N}_{B}$ must therefore match the set of the left-hand sides of the enabling relations: $\mathcal{N}_{B}=\{t|\exists x\in X,t\vdash x\}$.
    As a reminder, this yields the pattern shown in Figure 9. It also follows that $\textsf{enabled}(t)=\textsf{successor}(t)$.
  • Bob’s information sets
    $\mathcal{I}_{B}=\mathcal{N}_{B}$ necessarily because of BOB-S.
  • Bob-Alfred edges
    If we consider an information set $x$, the parent of each node in $x$ must be unique according to BA1. Furthermore, if we consider all enabling relations activating $x$, the parents of nodes in $x$ are exactly and all the values of $t$ such that $t\vdash x$ by definition of the enabling relation in Definition 7. It follows that edges from Bob to Alfred are of the form $(t,(t,x,c))$. Thus, $\mathcal{R}_{A}=\{(t,(t,x,c))|t\vdash x\}$ where the allowed values for $c$, which index the nodes in information set $x$ with parent $t$, $((t,x,c))_{c}$, must still be defined.
  • Bob-Alfred labels
    $\forall(t,(t,x,c))\in\mathcal{R}_{A},\sigma(t,(t,x,c))=c$ follows from the renaming convention of the labels from Bob to Alfred. According to this convention, for any node $t$ played by Bob, there are as many outgoing edges from $t$ as there are pairs $(x,d)$ such that $x\in d$, $d\in\chi(t)$, and $\sigma(t,(x,d))=d$. Thus $d$ also indexes the nodes in the information set $x$ and can be identified with $c$, and the edge going from $t$ to node $(t,x,c)$ must be labeled with $\sigma(t,(t,x,c))=c$.
  • Bob’s actions by node
    We now show that $\chi(t)=\mathcal{C}_{t}$. This amounts to showing that $\chi(t)=\{C\cap\textsf{successor}(t)|C\in\mathcal{C}\}\setminus\{\emptyset\}$.
    Because of the naming convention of Bob’s actions, we know that the labels in $\chi(t)$ are sets of information sets that are successors of $t$. Thus $\chi(t)\subseteq\mathcal{P}(\textsf{successor}(t))$,
    Recall that $\mathcal{C}$ is by construction of the causal contextuality scenario a natural cover. Let us take an element $c$ in $\mathcal{C}_{t}$. By definition, $c$ is not the empty set, and $c$ is the intersection of some context $C\in\mathcal{C}$ with the set of successors of $t$. Thus, there is a corresponding complete history $z$ of the original game such that $\forall i\in\textsf{successor}(t),i\in c\iff z(i)\neq\bot$. This means that $c$ contains exactly the information sets that are activated by $t$ in $z$, and because $c$ is not empty, there is at least one.
    Because of the naming convention of Bob’s nodes, it means that $c$ corresponds to the label born by the edges activating these information sets, and thus $c\in\chi(t)$.
    The converse is also true. If we take an action $c\in\chi(t)$, it labels the edges to all the nodes $(t,x,c)$ such that $x\in c$ because of the naming convention, and $c$ is not the empty set because of EVEN and BOB-A. Thus, there exists a complete history assigning $c$ to $t$ (otherwise $t$ would be an unused node, contradicting that the game is clean) and assigning each $x\in c$ to some value, and no value for the other successors of $t$. The active information sets of Alfred in this complete history form a set that is an element $C$ of $\mathcal{C}$, the natural cover, which is a superset of $c$. Thus $c=C\cap\textsf{successor}(t)$, and $c$ is not the empty set. Thus, $c\in\mathcal{C}_{t}$.
    We have thus proven that $\chi(t)=\mathcal{C}_{t}$.
    We also note that there cannot be any duplicates, i.e., there cannot be distinct nodes in $x$ with the same parent $t$ and indexed by the same $c\in\chi(t)$ because of BA2.
  • Alfred’s nodes
    We have shown that the nodes in an information set $x$ with parent $t$ (where $t\vdash x$) are indexed by $\chi(t)$, so that $\mathcal{N}_{A}=\{(t,x,c)|x\in X,t\vdash x,c\in\chi(t)\}$. Also, we have shown that $\chi(t)=\mathcal{C}_{t}$. Thus $\mathcal{N}_{A}=\{(t,x,c)|x\in X,t\vdash x,c\in\mathcal{C}_{t}\}$.
  • Alfred-Bob edges
    Since we have established the pattern in Figure 9, it follows that edges from Alfred to Bob connect each information set in the domain of $t$ to $t$. Thus, the edges from Alfred to Bob are given by $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{domain}(t)\}$. All these edges must be present (for all $u$ and $c$ such that $(u,x,c)$ is a node) because of AB1’s implying that the nodes in an information set must have the same successors.
  • Alfred-Bob labels
    Again since we have established the pattern in Figure 9, the label on an edge from Alfred’s information set $y$ to Bob’s node $t$ must be $t(y)$. It follows that $\forall((u,x,c),t)\in\mathcal{R}_{B},\sigma((u,x,c),t)=t(x)$
  • Bob’s actions
    $\mathcal{A}_{B}=\{\mathcal{C}_{t}|t\in\mathcal{N}_{B}\}$ because the game has no unused actions and we already showed that $\chi(t)=\mathcal{C}_{t}$.
  • Alfred’s actions by node
    $\forall(t,x,c)\in\mathcal{N}_{A},\chi(t,x,c)=O_{x}$. This is true by the construction of the causal contextuality scenario (second bullet point of Definition 7).
  • Alfred’s actions
    : $\mathcal{A}_{A}=\bigcup_{x}O_{x}$. This is true because there are no unused actions.
  • Assignment to Alfred
    We have $\forall n\in\mathcal{N}_{A},\rho(n)=\text{Alfred}$ by the construction of $\mathcal{N}_{A}$.
  • Assignment to Bob
    We have $\forall n\in\mathcal{N}_{B},\rho(n)=\text{Bob}$ by the construction of $\mathcal{N}_{B}$.
  • All nodes
    We must have $\mathcal{N}=\mathcal{N}_{A}\cup\mathcal{N}_{B}$ because there are only two players according to 2-PLAYERS.
  • All edges
    $\mathcal{R}=\mathcal{R}_{A}\cup\mathcal{R}_{B}$ because there are only two players according to 2-PLAYERS.
  • All actions
    $\mathcal{A}=\mathcal{A}_{A}\cup\mathcal{A}_{B}$ because there are only two players according to 2-PLAYERS.
  • All information sets
    $\mathcal{I}=\mathcal{I}_{A}\cup\mathcal{I}_{B}$ because there are only two players according to 2-PLAYERS.
5 Concrete impact of the mapping on the simplicity of constructs and proofs
We now turn to a few examples showing that looking at a causal contextuality scenario via its associated spacetime game is directly useful and simplifies certain constructs, reasonings, and proofs.
5.1 Good covers
In the paper introducing causal contextuality scenarios [8], it is conjectured that certain covers dubbed “good covers” characterize certain causal contextuality scenarios for which desirable properties apply—such as the sheaf property. We first show that these good covers are straightforward to define from the mapping provided in this paper; in the next section, we will show that it is also straightforward to prove that the sheaf property applies in the case of pure strategies.
Concretely, causal contextuality scenarios that have a good cover can be defined as the range of the mapping of Definition 7. Causal contextuality scenarios with a cover that is not a good cover can still be mapped to an alternating spacetime game, however, this mapping does not roundtrip and the correspondence has no semantic or physical meaning in this case.
Definition 9.
A causal contextuality scenario is said to have a good cover if it is in the range of the mapping defined in Definition 7.
Theorem 5.
The mapping of Definition 7 becomes bijective if we consider its corestriction to causal contextuality scenarios with good cover.
Proof.
Replacing the codomain of an injective function with the range of the function makes it bijective. ∎
It is also useful to know that one can change the cover of any causal contextuality scenario to be a good cover, keeping its other components intact.
Theorem 6.
Given a causal contextuality scenario, it is always possible to find a causal contextuality scenario with a good cover that only differs from the first one in the cover.
Proof.
Given a causal contextuality scenario, one can:
  • Replace its cover with the powerset of $X$.
  • Map it to an alternating spacetime game.
  • Map back to a causal contextuality scenario, which will have a good cover by construction.
Note that the powerset of $X$ is not necessarily a good cover, so that the final causal contextuality scenario’s good cover need not be the powerset of $X$. The suchly obtained causal contextuality scenario only differs from the initial causal contextuality scenario in the cover, as the mapping leaves the measurements, outcomes, and enabling relations intact. This is because (i) measurements correspond to Alfred’s information sets, left-hand sides of enabling relations to Bob’s nodes, and outcomes to Alfred’s actions at a given information set, (ii) since we start with the full powerset of $X$, no measurement, outcome or enabling relation is lost while mapping the causal contextuality scenario to a spacetime game (as could occur more generally with some covers that are not good covers), (iii) by definition, the mapping the spacetime game to the final causal contextuality scenario with good cover also does not lose any measurement, outcome or enabling relation. ∎
Note that the good cover is not generally unique given measurements, outcomes, and enabling relations. The determination of the most sensible cover to model a specific experiment requires additional knowledge about the considered Hilbert spaces, observables, and commutation properties.
5.2 Global strategies
The paper introducing causal contextuality scenarios [8] defines strategies of nature and of the experimenter for a given causal contextuality scenario. Follow-up work [20] also mentions that these strategies can be expressed as global strategies that map measurements to outcomes, a construct also referred to—in our understanding— as the “flattening construction” in an IQSA keynote speech.
If we view a causal contextuality scenario as its corresponding spacetime game, this game has a corresponding strategic form as we explained in 2020 [10]: it is the (reduced) strategic form of the associated game in extensive form with imperfect information, which was constructed 75 years ago by Harold Kuhn [12][9]. A pure strategy in this strategic form provides the desired global strategy function.
5.3 The sheaf of strategies
Let us take any causal contextuality scenario with good cover.
In the flat case, it was shown by Abramsky and Brandenburger [21] that the presheaf of strategies has the sheaf property because it also involves the gluing of partial functions on discrete space that agree on their overlaps. Thanks to the spacetime game framework, it is straightforward to translate this argument to any causal contextuality scenario with good cover.
Indeed, if we take a pure strategy as defined by Kuhn instead of the form given by Abramsky et al [8] in terms of game semantics, we note that these strategies are partial functions on a discrete space ($\mathcal{I}$) to actions. Thus, we can generally consider the restrictions of these partial functions to smaller domains, from which it follows as in shown by Abramsky and Brandenburger [21] that the suchly defined presheaf of strategies has the sheaf property.
This was left as an open question by [8] to determine “good covers” for which this is the case, and we have thus shown that Definition 9 successfully addresses this open question. The simplicity of the construction of the strategy sheaf of Nature demonstrates the expressive power of the spacetime game framework, itself building on 75 years of game theory literature.
5.4 Mixed strategies and empirical models
Nash equilibria [7] and other concepts apply naturally to spacetime games [10] through their associated extensive form. The strategic form constructed by Kuhn has pure strategies but also mixed strategies, which are probabilistic mixtures of pure strategies, i.e., a probability distribution over all pure strategies of a given player.
A mixed strategy of nature thus corresponds to a deterministic hidden variable model where $\lambda$ is a pure strategy. Free choice, as commonly assumed in the quantum foundations’ literature, translates to nature and the observers independently picking their mixed strategies – as is done when considering Nash equilibria in classical game theory.
Let us take the game in Figure 1 (Figure 2 for its extensive form with imperfect information) as an example. It corresponds to the 2-2-2 EPR experiment. If we observe that X and Y have disjoint supports (because Alice either picks basis x or y) and we merge them into a single one (call it X) – and we do the same for Bob (W and Z are merged into a single variable Z) – then the formula for computing the probability for the complete history resulting from a combination of a mixed strategy of nature (a distribution over the values of $\Lambda$) and the mixed strategies of observers (distributions over the values of $A$ and $B$) happens to exactly correspond to the formula used in the Bell inequality literature [22] for local-realist theories with measurement independence:
$P_{XZ|AB}=\sum_{\lambda}P_{\Lambda}(\lambda)P_{X|A\Lambda}P_{Z|B\Lambda}$
Figure 10 shows how this can also be interpreted as a causal network, with all variables depending deterministically on the mixed strategies, leading to a distribution of the complete histories for each combination of mixed strategies of Alice, Bob, and Alfred. Since in our non-Nashian paradigm [5][6] we posit a symmetry between nature and observers (or equivalently, between settings and outcomes), we add explicit lambdas also for the observers and not only for nature.
Figure 10: A visualization of the Nash mixed strategies as exogenous nodes in a causal network. We merged XY and WZ into single random variables, by observing that the supports are disjoint, to keep the visual simple to understand.
In the special case of flat games (no adaptive measurements), it was shown by Abramsky and Brandenburger [21] that the existence of a deterministic hidden variable model realizing an empirical model is implied by the existence of a global section for it. With spacetime games, this is intuitive to understand: we observe that a mixed strategy of nature is the same mathematical object as a global section of the event distribution presheaf and we remember that spacetime games have a natural cover given by the projection of the support of their complete histories to nature.
Then, the following result follows from Harold Kuhn’s strategic form:
Theorem 7.
Given a causal contextuality scenario with a good cover and an empirical model, the existence of a global section that is consistent with an empirical model implies the existence of a deterministic hidden variable model for this empirical model.
Proof.
If we consider the associated spacetime game, the existence of a global section consistent with the empirical model is equivalent to the existence of a mixed strategy of nature, i.e., a distribution on nature’s pure strategies, that is consistent with the empirical model. This mixed strategy provides a deterministic hidden variable model for this empirical model. ∎
We have just generalized the result of Abramsky and Brandenburger [21] from the flat case to any causal contextuality scenario with a good cover. The reason for the short proof is not that it is trivial; rather, it builds on previous work including Kuhn’s construct. This is consistent with our qualitative argument given by Baczyk and Fourny [6] that Nashian game theory is not compatible with quantum physics. The formulation in terms of presheaves and the sheaf property [21] strengthens this argument.
5.5 Perfect information
We observed [10] that imperfect information in the extensive form of the game corresponding to a spacetime game with perfect information can be interpreted in terms of the preservation of locality: information sets are grouped in the way they are because of the finite speed of light, and spacelike separation.
If we do not require perfect information for spacetime games, this introduces another physical justification for imperfect information in the resulting extensive form, namely, noncontextuality because non-singleton information sets in the spacetime game (which lead to the fusion of the corresponding information sets in the extensive form) arise from the different contexts in which a measurement can be carried out, with the result independent of the context. This is also consistent with the fact that locality is a special case of non-contextuality.
5.6 Temporal contextuality scenarios
Temporal contextuality scenarios were introduced by Searle et al [20] as a special case of causal contextuality scenario in which the measurements are timelike-separated. If we consider the associated spacetime games, temporal contextuality scenarios correspond to games in extensive form with perfect information. Furthermore, the lookbacks—corresponding to limited memory on past experiments— can be modeled in the spacetime game framework with imperfect information, i.e., by connecting identical measurements carried out in distinct pasts together. All the above simplified constructs apply: causal secured covers, global strategies, the presheaf of pure strategies, and the link between the existence of a global section of an empirical model and the existence of a deterministic hidden variable model.
6 Related work
Hardy models [14], PR boxes [15], and GHZ models [16] are representable with spacetime games.
The framework also maps to contextuality by default [23], which corresponds to having perfect information. Forcing random variables corresponding to measurement outcomes to have the same value formally corresponds to grouping the corresponding nodes into the same information set.
Sheaves of sections as constructed by Gogioso and Pinzani [13] correspond to a special case of spacetime games with perfect information in which measurements are performed independently of previous measurement outcomes, without merging the information sets (see also Section 6.b. in [8]), and with unique causal bridges. The fact that the sheaf property obtains is consistent with our results, which generalize the case covered by Gogioso and Pinzani [13] to all alternating spacetime games, and causal contextuality scenarios that have a natural cover.
A definition of free choice is given by Renner and Colbeck [24], as a decision (called spacetime variable) being independent of anything it could not have caused. A deterministic resolution algorithm for all spacetime games with perfect or imperfect information, and which circumvents impossibility theorems (such as the sheaf property of the strategy distribution presheaf not generally holding) by weakening free choice, is given in [5], with the special case of games in normal form [25] and of games in extensive form with perfect information [17]. A high-level discussion of the non-Nashian paradigm in the context of quantum physics interpreted as a game between the observer and nature is also available [26].
The seeding philosophy of the non-Nashian approach as well as the notion of causal bridge is attributable to Dupuy [18] (counterfactual decision theory) and the game theoretic seed and initial conjectures were made by [19].
There is also subsequent work by Kadnikov and Wichardt [27][28] that provides a strategic decision-making model generalizing the normal form and extensive form with graph structures related to spacetime games. This work defines the outcomes of the model based on sets of leaves of the graph structure and nicely connects it to the set-theoretic form by Bonanno [29][30]. These sets of leaves can be put in correspondence with the complete histories, and the outcomes of the game in extensive form with imperfect information, giving additional and complementary insights.
[31] give an excellent argument of why impossibility theorems do not disprove the existence of local theories. Hance and Hossenfelder [32] also are of the opinion that free choice can be weakened. Wharton and Argaman [33] refer to this possibility as (non-)future-input-dependent parameters, and it is also known as measurement independence (see for example the work of Storz et al [34] for the explicit mention of this assumption in Bell experiments) or setting independence [35].
7 Conclusion
We would like to conclude by sharing our scientific belief that only spacetime games with perfect information correspond to a physical reality, and that spacetime games with imperfect information are, above all, a useful mathematical tool for reductio ad absurdum proofs that nature is contextual. This is aligned with work by Dzhafarov [23] and on keeping locality as a fundamental principle of nature, provided we weaken measurement independence and assume non-Nashian free choice instead [6].
8 Acknowledgements
I thank my students at ETH who contributed to the game framework in various projects over the years: Felipe Sulser, Emilien Pilloud, Ramon Gomm, Luc Stoffer, Martin Kucera, Michael Baczyk, Florian Pauschitz.